OpenStudy (emilyjones284):

Find the coordinates of the vertex of the following equation: y equals 2x^2+8x-17

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) in your case \(x=2,b=8\) and \(-\frac{b}{2a}=-\frac{8}{2\times 2}=-2\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by \(-2\)